OpenStudy (anonymous):

Write a equation that will help you solve for d using the ratio sin40

OpenStudy (anonymous):



OpenStudy (phi):

do you know soh cah toa ?

OpenStudy (anonymous):

yeah i am supposed to use soh to figure this out so it would be opposite/hypotenuse

OpenStudy (anonymous):

unless my last question was wrong and the right answer is cah

OpenStudy (jdoe0001):

you are given the hypotenuse of "40", and also it's opposite, so

OpenStudy (anonymous):

so it is soh?

OpenStudy (jdoe0001):

yes

OpenStudy (phi):

opposite means you can draw an arrow from the angle to the other side: 

OpenStudy (phi):

notice the hypotenuse is opposite the right angle (the 90º angle)

OpenStudy (phi):

now you know to use soh sin(angle) = opposite/hypotenuse replace angle with 40, replace opposite with 200, replace hypotenuse with d

OpenStudy (anonymous):

how would i then solve that?

OpenStudy (phi):

first, what do you have ?

OpenStudy (anonymous):

sin (40) = 200/ d

OpenStudy (anonymous):

@phi

OpenStudy (phi):

yes. to solve, I like the "d" up top. one way to get the d up top is multiply both sides by d \[ d \cdot sin(40) = \frac{200}{d} \cdot d \] can you simplify the right hand side ?

OpenStudy (phi):

remember you can swap the order of the multiplies . keep the top 200 * d but sway d*1 to 1*d \[ \frac{200}{d} \cdot \frac{d}{1} =\frac{200}{1} \cdot \frac{d}{d} \] d/d is 1 , right ?

OpenStudy (phi):

and 200/1 * 1 is just 200

OpenStudy (phi):

you get \[ d \sin(40) = 200\] now divide both sides by sin(40) to get d by itself. to get a number, you will need to use a calculator.

OpenStudy (anonymous):

200 divided by 40 is =5

OpenStudy (anonymous):

and on the other side 40 will just cancel out right so what we will have is d=5? @phi

OpenStudy (phi):

right idea, but sin(40) is not 40. get a calculator or use google: type this into the search window sin(40 degrees)=

OpenStudy (anonymous):

how would i type this into my calculator

OpenStudy (phi):

try sin 40 enter

OpenStudy (phi):

or sin 40 =

OpenStudy (anonymous):

ok i did

OpenStudy (phi):

what did you get ?

OpenStudy (phi):

d= 200/sin(40)

OpenStudy (anonymous):

i got .64278

OpenStudy (phi):

yes, and d= 200/ugly number

OpenStudy (anonymous):

312.5

OpenStudy (phi):

I am getting something slightly smaller. don't round type in 200 / sin 40 =

OpenStudy (anonymous):

oh ok i got 311.144

OpenStudy (phi):

yes. as a check, if we do inverse sin( 200/311.144) we should get back 40 degrees (or close to it)